OpenStudy (anonymous):

what 2 consecutive numbers add to 257

OpenStudy (anonymous):

128 and 129

jimthompson5910 (jim_thompson5910):

x+(x+1)=257 2x+1=257 2x=257-1 2x=256 x=256/2 x=128 So the two numbers are 128 and 129